OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. R(0, _), S(-4, 2) m = 5

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

\[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

Hello there, Welcome to OpenStudy! @kristyb056

OpenStudy (anonymous):

The slope formula is \(\sf Slope=\Large\frac{y_2-y_1}{x_2-x_1}\) Where \(\sf (x_1,y_1), (x_2,y_2)\) are the two points.

OpenStudy (anonymous):

\[5=\frac{ 2-y_1 }{ -4-0 }\]

OpenStudy (anonymous):

\[5=\frac{ 2-y_1 }{ -4 }\]

OpenStudy (anonymous):

so jow do i find y1?

OpenStudy (anonymous):

how*

OpenStudy (anonymous):

first,multiply -4 on both side

OpenStudy (anonymous):

So you are given two points, \(\sf (0, _-), (-4,2)\) And so then \(\sf x_1 = 0 \\ y_1= _- \\ x_2 = -4 \\ y_2=2\) And we are looking for \(\sf y_1\)

OpenStudy (anonymous):

\[5(-4)=-4(\frac{ 2-y_1 }{ -4 })\]

OpenStudy (anonymous):

I will let @MARC_ continue on helping you as I do not want to confuse you with two people talking at the same time.

OpenStudy (anonymous):

\[-20=2-y_1\]

OpenStudy (anonymous):

so would i subtract both sides?

OpenStudy (anonymous):

add?

OpenStudy (anonymous):

\[y_1=2+20\]

OpenStudy (anonymous):

y1=22?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

So thats the answer? Thank you!

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Find the slope of the line that contains the points named. X(2, -1), Y(0, 4)

OpenStudy (anonymous):

would it be (2-0)(-1-4)

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

substitute the given values into the equation

OpenStudy (anonymous):

(0+1)/(4-2)

OpenStudy (anonymous):

(2,-1)(0,4)\[(x_1,y_1)(x_2,y_2)\]

OpenStudy (anonymous):

ugh i keep getting that missed up

OpenStudy (anonymous):

\[m=\frac{ y_2-y_1 }{ x_2-x_1 }\] plug in the given values into this equation

OpenStudy (anonymous):

\[=\frac{ 4-(-1) }{ 0-2 }\]

OpenStudy (anonymous):

2.5?

OpenStudy (anonymous):

you can just leave the answer in fraction form

OpenStudy (anonymous):

Find the slope of the line that contains the points named. J(1, -4), K(3, -1)

OpenStudy (anonymous):

(3-1) (-1-4)